# Optimal Computation in Anonymous Dynamic Networks

### v7

@@ 3.1 Definition of History Tree
Figure 1: The first rounds of a dynamic network with $n=9$ processes and the corresponding levels of the history tree. Level $L_{t}$ consists of all nodes at distance $t+1$ from the root $r$, which represent indistinguishable processes at the end of the $t$th communication round. The multiplicities of the red multi-edges of the history tree are explicitly indicated only when greater than $1$. The letters A, B, C denote processes’ inputs; all other labels have been added for the reader’s convenience, and indicate classes of indistinguishable processes (non-trivial classes are also indicated by dashed blue lines in the network). Note that the two processes in $b_{4}$ are still indistinguishable at the end of round $2$, although they are linked to the distinguishable processes $b_{5}$ and $b_{6}$. This is because such processes were in the same class $a_{5}$ at round $1$. The subgraph induced by the vertices in the green blob is the view of the two processes in $b_{1}$. None of the levels of this view is complete, except $L_{-1}$.
@@
Indistinguishable processes. In an anonymous network, processes can only be distinguished by their inputs or by the multisets of messages they have received thus far. This leads to an inductive definition of indistinguishability: Two processes are indistinguishable at the end of round 00 if and only if they have the same input. At the end of round $t\geq 1$, two processes $p$ and $q$ are indistinguishable if and only if they were indistinguishable at the end of round $t-1$ and, for every equivalence class $A$ of processes that were indistinguishable at the end of round $t-1$, both $p$ and $q$ receive an equal number of (identical) messages from processes in $A$ at round $t$.
@@
Levels of a history tree. A history tree is a structure associated with a dynamic network. It is an infinite graph whose nodes are subdivided into levels $L_{-1}$, $L_{0}$, $L_{1}$, $L_{2}$, …, where each node in layer $L_{t}$, with $t\geq 0$, represents an equivalence class of processes that are indistinguishable at the end of round $t$. The level $L_{-1}$ contains a unique node $r$, representing all processes in the system.
@@
In addition, each node of $L_{0}$ has a label denoting the input of the processes it represents (by definition of indistinguishability at round 00, each node of $L_{0}$ represents all processes with a certain input). Nodes not in $L_{0}$ do not have any labels.
@@
Black and red edges. A history tree has two types of undirected edges; each edge connects nodes in consecutive levels. The black edges induce an infinite tree rooted at $r$ and spanning all nodes. A black edge $\{v,v^{\prime}\}$, with $v\in L_{t}$ and $v^{\prime}\in L_{t+1}$, indicates that the child node $v^{\prime}$ represents a subset of the processes represented by the parent node $v$.
@@
The red multiedges represent messages. Each red edge $\{v,v^{\prime}\}$ with multiplicity $m$, with $v\in L_{t}$ and $v^{\prime}\in L_{t+1}$, indicates that, at round $t+1$, each process represented by $v^{\prime}$ receives a total of $m$ (identical) messages from processes represented by $v$.
@@
Anonymity of a node. The anonymity $a(v)$ of a node $v$ of a history tree is defined as the number of processes represented by $v$. Since the nodes in a level represent a partition of all the processes, the sum of their anonymities must be equal to the total number of processes in the system, $n$. Moreover, by the definition of black edges, the anonymity of a node is equal to the sum of the anonymities of its children.
@@
View of a process. A monotonic path in the history tree is a sequence of nodes in distinct levels, such that any two consecutive nodes are connected by a black or a red edge. For any process $p$, let $h(p,t)$ be the unique node representing $p$ in the level $L_{t}$ of the history tree. We define the view of $p$ at round $t$ as the finite subgraph of the history tree induced by all the nodes spanned by monotonic paths with endpoints $h(p,t)$ and the root $r$.
A node of the history tree is missing from a view if it is not among the nodes in the view. A level of a view is complete if no node is missing from that level. Observe that, if a level of a view is complete, then all previous levels of the view are also complete.
3.2 Construction of Views
Intuitively, the view of a process represents the portion of the history tree that the process “knows” at that time. In fact, there is an effective procedure that allows all processes in a network to locally construct a representation of their view at all rounds. We will present this procedure and discuss some related results, including the fundamental theorem of history trees, Theorem 3.1.
Merge operation. The basic operation that is used to construct and update views is called merge, and is illustrated in Figure 2. Merging two views $\mathcal{V}$ and $\mathcal{V}^{\prime}$ is a very natural operation whose result is the minimum graph that contains both $\mathcal{V}$ and $\mathcal{V}^{\prime}$ as induced subgraphs.
@@ 3.2 Construction of Views
Figure 2: Updating the view of a process represented by node $v$ after it receives messages from a process represented by $u$ and a process represented by $w$. The node $v^{\prime}$ is created as a child of $v$, the views corresponding to $u$ and $w$ are merged with the old view, and red edges are created connecting $v^{\prime}$ with $u$ and $w$.
Local construction. We will now describe a local algorithm that allows processes to construct and update their views at every round. That is, assuming that each process has its view as its internal state at the beginning of round $t$ and sends its view to all its neighbors at round $t$, there is a local algorithm $\mathcal{A}^{\ast}$ that allows the process to construct its new view at the end of round $t$.
Constructing the view at round 00 is simple, because the process just has to read its own input and create a root $r$ with a single child bearing a label equal to its input.
After that, the view can be updated at every round as shown in Figure 2. Namely, the node $v$ that currently represents the process gets a new child $v^{\prime}$, representing the same process in the next round. Then, if $m\geq 1$ copies of a view $\mathcal{V}$ are received as messages, then $\mathcal{V}$ is merged into the current view, and a red edge with multiplicity $m$ is added to connect $v^{\prime}$ with the deepest node of $\mathcal{V}$. The correctness of this operation follows straightforwardly from the definition of view.
Size of the view. We remark that the size of the view of a process at round $t$ is polynomial in $t$ and $n$, and so is the amount of local computation needed to update the process’ view at round $t$ based on the messages it has received. More specifically, the view can be represented in $O(tn^{2}\log M)$ bits, where $M$ is the maximum number of messages that any process may receive at any round (if the network is simple, then $M<n$).
Indeed, each of the $O(t)$ levels in the view contains up to $n$ nodes. Thus, there are $O(n^{2})$ possible red edges between any two consecutive levels, each of which has a multiplicity that can be represented in $O(\log M)$ bits.
Fundamental theorem. We will now give a concrete meaning to the idea that the view of a process at round $t$ contains all the information that the process can use at that round. This intuition is made precise by the following fundamental theorem of history trees.
@@
If all processes in a system execute the same local algorithm $\mathcal{A}$ at every round, then the internal state of a process $p$ at the end of round $t$ is determined by a function $\mathcal{F}_{\mathcal{A}}$ of the view of $p$ at round $t$. The mapping $\mathcal{F}_{\mathcal{A}}$ depends entirely on the algorithm $\mathcal{A}$ and is independent of $p$.
@@
Denote by $h(p,t)$ the node of the history tree that represents $p$ at round $t$, by $\mathcal{V}(p,t)$ the view of $p$ at round $t$, and by $\sigma(p,t)$ the internal state of $p$ at round $t$. The construction of $\mathcal{F}_{\mathcal{A}}$ is done by induction on $t$ in such a way that $\mathcal{F}_{\mathcal{A}}(\mathcal{V}(p,t))=\sigma(p,t)$ for any process $p$ and any round $t$.
If $t=0$, then $\mathcal{V}(p,t)$ consists of a single node $h(p,t)$ in $L_{0}$ attached to the root $r$. Since each node of $L_{0}$ is labeled as the input of the processes it represents, $\mathcal{F}_{\mathcal{A}}$ can extract the input of $p$ from $h(p,t)$ and use it to compute its internal state $\sigma(p,t)$, which at round 00 is just a function of the input.
Let $t>0$ and assume that the inductive hypothesis $\mathcal{F}_{\mathcal{A}}(\mathcal{V}(p,t-1))=\sigma(q,t-1)$ holds for any process $q$. Let $p$ be any process, and let us prove that $\mathcal{F}_{\mathcal{A}}(\mathcal{V}(p,t))=\sigma(p,t)$. Observe that, by definition, if a view contains a node in level $L_{i}$, it also contains the view of the processes represented by that node at round $i$.
It is easy to infer $h(p,t)$ from $\mathcal{V}(p,t)$, because it is the unique node of maximum depth. Thus, the node $h(p,t-1)$ is the parent of $h(p,t)$, which is in $\mathcal{V}(p,t)$ because it is connected to $h(p,t)$ by a black edge. It follows that $\mathcal{F}_{\mathcal{A}}$ can extract from $\mathcal{V}(p,t)$ the view of $p$ at round $t-1$ and, by the inductive hypothesis, it can compute $\sigma(p,t-1)$.
Also, for each process $q$ that sends messages to $p$ at round $t$, the node $h(q,t-1)$ is in $\mathcal{V}(p,t)$, because it is connected to $h(p,t)$ by a red edge with a certain multiplicity $m\geq 1$. Hence, $\mathcal{F}_{\mathcal{A}}$ can extract $m$ from $\mathcal{V}(p,t)$, as well as the view of $q$ at round $t-1$, and compute $\sigma(q,t-1)$ by the inductive hypothesis. From the internal state $\sigma(q,t-1)$, the message received by $p$ from $q$ at round $t$ can also be computed.
@@
As a consequence, all the processes represented by the same node in level $L_{t}$ of the history tree must have the same state (and give the same output) at the end of round $t$, regardless of the deterministic algorithm being executed. This is in agreement with the idea that the processes represented by the same node are indistinguishable.
@@
At the end of any round $t$, all processes represented by the same node in $L_{t}$ have the same internal state.
@@
If two processes are represented by the same node of $L_{t}$, they have the same view at round $t$. Thus, by Theorem 3.1, they have the same internal state at the end of round $t$. ∎
Significance. The significance of the fundamental theorem is that it allows us to shift our focus from dynamic networks to history trees. Recall that there is a local algorithm $\mathcal{A}^{\ast}$ that allows processes to construct and update their view at every round. Now, Theorem 3.1 guarantees that processes do not lose any information if they simply execute $\mathcal{A}^{\ast}$, regardless of their goal, and then compute their task-dependent outputs as a function of their respective views. Thus, in the following, we will assume without loss of generality that the internal state of every process at every round, as well as all the messages it sends to all its neighbors, always coincide with its view at that round.
@@ 3.3 Related Concepts
Views of Yamashita–Kameda. Precise graph-theoretic characterizations of when certain fundamental problems for anonymous static networks are solvable were later given by Yamashita and Kameda [57]. The same authors also introduced the concept of view of a process in a static network whose links are endowed with port numbers (i.e., the links incident to the same process have distinct IDs) [56]. For Yamashita and Kameda, the view of a process $p$ in a static network $G$ is an infinite tree rooted at $p$ that encodes all the information about the network that can possibly be gathered by $p$. In the language of topological graph theory, the view of $p$ is akin to the universal cover of $G$ (i.e., the “largest possible” cover of $G$) rooted at $p$.
The notion of “view” in the sense of Yamashita–Kameda should not be confused with the views of history trees introduced in Section 3.1.191919We emphasize that, outside of Section 3.3, we will use the word “view” only in reference to history trees. However, the two concepts are loosely related, as one may construct the view of a process in a history tree at round $t$ by inspecting the Yamashita–Kameda view of this process truncated at depth $t$, and vice versa.
If processes with isomorphic views are considered identical, one obtains the quotient graph in the sense of Yamashita–Kameda. In the language of topological graph theory, this can also be characterized as the “smallest possible” graph that is covered by $G$.
@@
In summary, previous structures and theories related to anonymous static networks can effectively be reinterpreted within the framework of history trees. The advantage of using history trees is that they also incorporate information about the time at which specific classes of processes become distinguishable. This “temporal dimension” makes history trees an ideal tool for dynamic networks, where one cannot rely on topological regularity to infer temporal information.
Algorithms for static networks. From a computational standpoint, it is important to consider how a process in a network can algorithmically construct the aforementioned data structures and use them effectively to solve certain problems.
We already pointed out that the views of Yamashita–Kameda contain all the information that can be gathered locally by the processes; the same can be said of the minimum bases of Boldi–Vigna. Hence, constructing these structures allows a process to access all the information that it can possibly use for computations.
Yamashita and Kameda showed that, at round $t$, any process can algorithmically construct its view truncated at depth $t$. Thus, it can incrementally construct its view level by level. Moreover, they showed that, in a static network of $n$ processes, if two processes have isomorphic views up to depth $n^{2}$, they have isomorphic views at any depth [56]. This leads to the conclusion that, if the number $n$ is known to the processes, they can effectively solve general problems and terminate in $O(n^{2})$ communication rounds.
For example, since processes with isomorphic views are always indistinguishable, it is possible to elect a unique leader in a static network in a finite number of rounds if and only if there is a unique process whose view truncated at depth $n^{2}$ is distinct from all others.
Linear bounds. The $n^{2}$ upper bound was later improved by Norris, who showed that truncating views at depth $n-1$ is sufficient and may be necessary in some networks. That is, if two processes have isomorphic views up to depth $n-1$, they have isomorphic views at any depth [46].
@@
Proving this statement is immediate if one reframes it in terms of history trees. We have already argued that, in the history tree of a static network, the number of nodes in a level increases by at least one unit per level until level $L_{s}$, where it becomes maximum. That is, $|L_{s}|=\widehat{n}$, and clearly $s<\widehat{n}$, because $|L_{0}|\geq 1$. Thus, if two processes have isomorphic views (in the history tree) at round $\widehat{n}-1$, they have isomorphic views at all rounds, which is equivalent to the statement of Fraigniaud and Pelc.
Size of a view. When designing efficient algorithms, one also wishes to optimize memory usage and message sizes, as well as running times. Unfortunately, in the worst case, a view of Yamashita and Kameda truncated at depth $t$ grows exponentially in $t$, making their approach unsuitable for systems where memory limitations are an issue.
For this reason, Tani proposed a method to efficiently compress a truncated view by identifying isomorphic subtrees within the view itself. Using Tani’s technique, a compressed view (of an undirected network with port numbers) truncated at depth $t$ can be stored in $O(tnM\log M)$ bits, where $M$ is the maximum degree of a process in the network [54].
This essentially matches the upper bound on the size of a view of a history tree of a $\tau$-union-connected network at round $t$, which is $O(tn^{2}\log\tau M)$ bits (the small discrepancy is due to the fact that our networks are modeled by disconnected dynamic multigraphs, as opposed to connected static simple graphs). Thus, for instance, our terminating Counting algorithm based on history trees requires $O(n^{3}\log\tau M)$ bits of memory per process in the worst case.
@@
A similar notion of causal influence can be formulated in the framework of history trees by stating that a process $p$ exerts a causal influence on a process $q$ at round $t$ if the view of $q$ at round $t$ includes a node $v$ representing $p$. It is important to note that not all processes represented by $v$ necessarily initiate a sequence of messages reaching $q$ by round $t$. However, at least one process represented by $v$ does so, and the specific identity of this process is irrelevant, since processes are anonymous.
@@
Kuhn et al. established that, in a 1-union-connected dynamic network, the set of processes that are causally influenced by a given process grows by at least one unit at every round, until all $n$ processes have been influenced. We can rephrase this observation in the language of history trees as follows.
  • The view of any process at any round $t$ contains nodes representing at least $\min\{t+1,n\}$ processes.
  • For any process $p$ and any round $t$, there are at least $\min\{t+1,n\}$ processes whose view at round $t$ contains a node representing $p$.
Broadcasting speed. As a consequence of the above, broadcasting a message, i.e., forwarding a message from a single process to the whole dynamic network, takes at most $n-1$ rounds. Therefore, the dynamic diameter $d$, defined as the maximum number of rounds it takes to forward a message from a process to any other process at any point in time, satisfies $d\leq n-1$ in any 1-union-connected dynamic network.
@@
Let $P$ be a set of processes in a 1-union-connected dynamic network of size $n$, such that $11\leq|P|\leq n-$ 1, and let $t\geq 0$. Then, at every round $t^{\prime}\geq t+|P|$, in the history of every process there is a node at level $L_{t}$ representing at least one process not in $P$.
@@
Let $Q$ be the complement of $P$ (note that $Q$ is not empty), and let $Q_{t+i}$ be the set of processes represented by the nodes in $L_{t+i}$ whose view contains a node in $L_{t}$ representing at least one process in $Q$. We will prove by induction that $|Q_{t+i}|\geq|Q|+i$ for all $0\leq i\leq|P|$. The base case holds because $Q_{t}=Q$. The induction step is implied by $Q_{t+i}\subsetneq Q_{t+i+1}$, which holds for all $0\leq i<|P|$ as long as $|Q_{t+i}|<n$. Indeed, because $G_{t+i+1}$ is connected, it must contain a link between a process $p\in Q_{t+i}$ and a process $q\notin Q_{t+i}$. Thus, the history of $q$ at round $t+i+1$ contains the history of $p$ at round $t+i$, and so $Q_{t+i}\subsetneq Q_{t+i}\cup\{q\}\subseteq Q_{t+i+1}$.
Now, plugging $i:=|P|$, we get $|Q_{t+|P|}|=n$. In other words, the history of each node in $L_{t+|P|}$ (and hence in subsequent levels) contains a node at level $L_{t}$ representing a process in $Q$. ∎

### v8

@@ 3.1 Definition of History Tree
Figure 1: The first rounds of a dynamic network with $n=9$ agents and the corresponding levels of the history tree. Level $L_{t}$ consists of all nodes at distance $t+1$ from the root $r$, which represent indistinguishable agents at the end of the $t$th communication round. The multiplicities of the red multi-edges of the history tree are explicitly indicated only when greater than $1$. The letters A, B, C denote agents’ inputs; all other labels have been added for the reader’s convenience, and indicate classes of indistinguishable agents (non-trivial classes are also indicated by dashed blue lines in the network). Note that the two agents in $b_{4}$ are still indistinguishable at the end of round $2$, although they are linked to the distinguishable agents $b_{5}$ and $b_{6}$. This is because such agents were in the same class $a_{5}$ at round $1$. The subgraph induced by the vertices in the green blob is the vista of the two agents in $b_{1}$. None of the levels of this vista is complete, except $L_{-1}$.
@@
Indistinguishable agents. In an anonymous network, agents can only be distinguished by their inputs or by the multisets of messages they have received thus far. This leads to an inductive definition of indistinguishability: two agents are indistinguishable at the end of round 00 if and only if they have the same input. At the end of round $t\geq 1$, two agents $p$ and $q$ are indistinguishable if and only if they were indistinguishable at the end of round $t-1$ and, for every equivalence class $A$ of agents that were indistinguishable at the end of round $t-1$, both $p$ and $q$ receive an equal number of (identical) messages from agents in $A$ at round $t$.
@@
Levels of a history tree. A history tree is a structure associated with a dynamic network. It is an infinite graph whose nodes are subdivided into levels $L_{-1}$, $L_{0}$, $L_{1}$, $L_{2}$, …, where each node in layer $L_{t}$, with $t\geq 0$, represents an equivalence class of agents that are indistinguishable at the end of round $t$. The level $L_{-1}$ contains a unique node $r$, representing all agents in the system.
@@
In addition, each node of $L_{0}$ has a label denoting the input of the agents it represents (by definition of indistinguishability at round 00, each node of $L_{0}$ represents all agents with a certain input). Nodes not in $L_{0}$ do not have any labels.
@@
Black and red edges. A history tree has two types of undirected edges; each edge connects nodes in consecutive levels. The black edges induce an infinite tree rooted at $r$ and spanning all nodes. A black edge $\{v,v^{\prime}\}$, with $v\in L_{t}$ and $v^{\prime}\in L_{t+1}$, indicates that the child node $v^{\prime}$ represents a subset of the agents represented by the parent node $v$.
@@
The red multiedges represent messages. Each red edge $\{v,v^{\prime}\}$ with multiplicity $m$, with $v\in L_{t}$ and $v^{\prime}\in L_{t+1}$, indicates that, at round $t+1$, each agent represented by $v^{\prime}$ receives a total of $m$ (identical) messages from agents represented by $v$.
@@
Anonymity of a node. The anonymity $a(v)$ of a node $v$ of a history tree is defined as the number of agents represented by $v$. Since the nodes in a level represent a partition of all the agents, the sum of their anonymities must be equal to the total number of agents in the system, $n$. Moreover, by the definition of black edges, the anonymity of a node is equal to the sum of the anonymities of its children.
@@
Vista of an agent. A monotonic path in the history tree is a sequence of nodes in distinct levels, such that any two consecutive nodes are connected by a black or a red edge. For any agent $p$, let $h(p,t)$ be the unique node representing $p$ in the level $L_{t}$ of the history tree. We define the vista [^2] The preliminary versions of this paper [26, 27] used the term “view” instead of “vista”. In the present version, we have adopted the new terminology to avoid confusion with the loosely related concept of view introduced by Yamashita and Kameda in the context of static networks [66]. The relationship between vistas and views is discussed in Section 3.3 and further explored in [65]. of $p$ at round $t$ as the finite subgraph of the history tree induced by all the nodes spanned by monotonic paths with endpoints $h(p,t)$ and the root $r$.
A node of the history tree is missing from a vista if it is not among the nodes in the vista. A level of a vista is complete if no node is missing from that level. Observe that, if a level of a vista is complete, then all previous levels of the vista are also complete.
3.2 Construction of Vistas
Intuitively, the vista of an agent represents the portion of the history tree that the agent “knows” at that time. In fact, there is an effective procedure that allows all agents in a network to locally construct a representation of their vista at all rounds. We will present this procedure and discuss some related results, including the fundamental theorem of history trees, Theorem 3.1.
Merge operation. The basic operation that is used to construct and update vistas is called merge, and is illustrated in Figure 2. Merging two vistas $\mathcal{V}$ and $\mathcal{V}^{\prime}$ is a very natural operation whose result is the minimum graph that contains both $\mathcal{V}$ and $\mathcal{V}^{\prime}$ as induced subgraphs.
@@ 3.2 Construction of Vistas
Figure 2: Updating the vista of an agent represented by node $v$ after it receives messages from an agent represented by $u$ and an agent represented by $w$. The node $v^{\prime}$ is created as a child of $v$, the vistas corresponding to $u$ and $w$ are merged with the old vista, and red edges are created connecting $v^{\prime}$ with $u$ and $w$.
Local construction. We will now describe a local algorithm that allows agents to construct and update their vistas at every round. That is, assuming that each agent has its vista as its internal state at the beginning of round $t$ and sends its vista to all its neighbors at round $t$, there is a local algorithm $\mathcal{A}^{\ast}$ that allows the agent to construct its new vista at the end of round $t$.
Constructing the vista at round 00 is simple, because the agent just has to read its own input and create a root $r$ with a single child bearing a label equal to its input.
After that, the vista can be updated at every round as shown in Figure 2. Namely, the node $v$ that currently represents the agent gets a new child $v^{\prime}$, representing the same agent in the next round. Then, if $m\geq 1$ copies of a vista $\mathcal{V}$ are received as messages, then $\mathcal{V}$ is merged into the current vista, and a red edge with multiplicity $m$ is added to connect $v^{\prime}$ with the deepest node of $\mathcal{V}$. The correctness of this operation follows straightforwardly from the definition of vista.
Size of the vista. We remark that the size of the vista of an agent at round $t$ is polynomial in $t$ and $n$, and so is the amount of local computation needed to update the agent’s vista at round $t$ based on the messages it has received. More specifically, the vista can be represented in $O(tn^{2}\log M)$ bits, where $M$ is the maximum number of messages that any agent may receive at any round (if the network is simple, then $M<n$).
Indeed, each of the $O(t)$ levels in the vista contains up to $n$ nodes. Thus, there are $O(n^{2})$ possible red edges between any two consecutive levels, each of which has a multiplicity that can be represented in $O(\log M)$ bits.
Fundamental theorem. We will now give a concrete meaning to the idea that the vista of an agent at round $t$ contains all the information that the agent can use at that round. This intuition is made precise by the following fundamental theorem of history trees.
@@
If all agents in a system execute the same local algorithm $\mathcal{A}$ at every round, then the internal state of an agent $p$ at the end of round $t$ is determined by a function $\mathcal{F}_{\mathcal{A}}$ of the vista of $p$ at round $t$. The mapping $\mathcal{F}_{\mathcal{A}}$ depends entirely on the algorithm $\mathcal{A}$ and is independent of $p$.
@@
Denote by $h(p,t)$ the node of the history tree that represents $p$ at round $t$, by $\mathcal{V}(p,t)$ the vista of $p$ at round $t$, and by $\sigma(p,t)$ the internal state of $p$ at round $t$. The construction of $\mathcal{F}_{\mathcal{A}}$ is done by induction on $t$ in such a way that $\mathcal{F}_{\mathcal{A}}(\mathcal{V}(p,t))=\sigma(p,t)$ for any agent $p$ and any round $t$.
If $t=0$, then $\mathcal{V}(p,t)$ consists of a single node $h(p,t)$ in $L_{0}$ attached to the root $r$. Since each node of $L_{0}$ is labeled as the input of the agents it represents, $\mathcal{F}_{\mathcal{A}}$ can extract the input of $p$ from $h(p,t)$ and use it to compute its internal state $\sigma(p,t)$, which at round 00 is just a function of the input.
Let $t>0$ and assume that the inductive hypothesis $\mathcal{F}_{\mathcal{A}}(\mathcal{V}(p,t-1))=\sigma(q,t-1)$ holds for any agent $q$. Let $p$ be any agent, and let us prove that $\mathcal{F}_{\mathcal{A}}(\mathcal{V}(p,t))=\sigma(p,t)$. Observe that, by definition, if a vista contains a node in level $L_{i}$, it also contains the vista of the agents represented by that node at round $i$.
It is easy to infer $h(p,t)$ from $\mathcal{V}(p,t)$, because it is the unique node of maximum depth. Thus, the node $h(p,t-1)$ is the parent of $h(p,t)$, which is in $\mathcal{V}(p,t)$ because it is connected to $h(p,t)$ by a black edge. It follows that $\mathcal{F}_{\mathcal{A}}$ can extract from $\mathcal{V}(p,t)$ the vista of $p$ at round $t-1$ and, by the inductive hypothesis, it can compute $\sigma(p,t-1)$.
Also, for each agent $q$ that sends messages to $p$ at round $t$, the node $h(q,t-1)$ is in $\mathcal{V}(p,t)$, because it is connected to $h(p,t)$ by a red edge with a certain multiplicity $m\geq 1$. Hence, $\mathcal{F}_{\mathcal{A}}$ can extract $m$ from $\mathcal{V}(p,t)$, as well as the vista of $q$ at round $t-1$, and compute $\sigma(q,t-1)$ by the inductive hypothesis. From the internal state $\sigma(q,t-1)$, the message received by $p$ from $q$ at round $t$ can also be computed.
@@
As a consequence, all the agents represented by the same node in level $L_{t}$ of the history tree must have the same state (and give the same output) at the end of round $t$, regardless of the deterministic algorithm being executed. This is in agreement with the idea that the agents represented by the same node are indistinguishable.
@@
At the end of any round $t$, all agents represented by the same node in $L_{t}$ have the same internal state.
@@
If two agents are represented by the same node of $L_{t}$, they have the same vista at round $t$. Thus, by Theorem 3.1, they have the same internal state at the end of round $t$. ∎
Significance. The significance of the fundamental theorem is that it allows us to shift our focus from dynamic networks to history trees. Recall that there is a local algorithm $\mathcal{A}^{\ast}$ that allows agents to construct and update their vista at every round. Now, Theorem 3.1 guarantees that agents do not lose any information if they simply execute $\mathcal{A}^{\ast}$, regardless of their goal, and then compute their task-dependent outputs as a function of their respective vistas. Thus, in the following, we will assume without loss of generality that the internal state of every agent at every round, as well as all the messages it sends to all its neighbors, always coincide with its vista at that round.
@@ 3.3 Related Concepts
Views of Yamashita–Kameda. Precise graph-theoretic characterizations of when certain fundamental problems for anonymous static networks are solvable were later given by Yamashita and Kameda [67]. The same authors also introduced the concept of view of an agent in a static network whose links are endowed with port numbers (i.e., the links incident to the same agent have distinct IDs) [66]. For Yamashita and Kameda, the view of an agent $p$ in a static network $G$ is an infinite tree rooted at $p$ that encodes all the information about the network that can possibly be gathered by $p$. In the language of topological graph theory, the view of $p$ is akin to the universal cover of $G$ (i.e., the “largest possible” cover of $G$) rooted at $p$.
The views introduced by Yamashita and Kameda are loosely related to the vistas of history trees defined in Section 3.1, and their exact relationship has been detailed in [65]. Notably, in static networks, one can construct an agent’s vista at round $t$ by inspecting the Yamashita–Kameda view of that agent truncated at depth $t$, and vice versa. We stress that this correspondence holds only in static networks, as the Yamashita–Kameda views are not defined for dynamic networks.
If agents with isomorphic views are considered identical, one obtains the quotient graph in the sense of Yamashita–Kameda. In the language of topological graph theory, this can also be characterized as the “smallest possible” graph that is covered by $G$.
@@
In summary, previous structures and theories related to anonymous static networks can effectively be reinterpreted within the framework of history trees. The advantage of using history trees is that they also incorporate information about the time at which specific classes of agents become distinguishable. This “temporal dimension” makes history trees an ideal tool for dynamic networks, where one cannot rely on topological regularity to infer temporal information.
Algorithms for static networks. From a computational standpoint, it is important to consider how an agent in a network can algorithmically construct the aforementioned data structures and use them effectively to solve certain problems.
We already pointed out that the views of Yamashita–Kameda contain all the information that can be gathered locally by the agents; the same can be said of the minimum bases of Boldi–Vigna. Hence, constructing these structures allows an agent to access all the information that it can possibly use for computations.
Yamashita and Kameda showed that, at round $t$, any agent can algorithmically construct its view truncated at depth $t$. Thus, it can incrementally construct its view level by level. Moreover, they showed that, in a static network of $n$ agents, if two agents have isomorphic views up to depth $n^{2}$, they have isomorphic views at any depth [66]. This leads to the conclusion that, if the number $n$ is known to the agents, they can effectively solve general problems and terminate in $O(n^{2})$ communication rounds.
For example, since agents with isomorphic views are always indistinguishable, it is possible to elect a unique leader in a static network in a finite number of rounds if and only if there is a unique agent whose view truncated at depth $n^{2}$ is distinct from all others.
Linear bounds. The $n^{2}$ upper bound was later improved by Norris, who showed that truncating views at depth $n-1$ is sufficient and may be necessary in some networks. That is, if two agents have isomorphic views up to depth $n-1$, they have isomorphic views at any depth [54].
@@
Proving these statements is immediate if one reframes them in terms of history trees. We have already argued that, in the history tree of a static network, the number of nodes in a level increases by at least one unit per level until level $L_{s}$, where it becomes maximum. That is, $|L_{s}|=\widehat{n}$, and clearly $s<\widehat{n}$, because $|L_{0}|\geq 1$. Thus, if two agents have isomorphic vistas (in the history tree) at round $\widehat{n}-1$, they have isomorphic vistas at all rounds, which is equivalent to the statement of Fraigniaud and Pelc.
Size of a view. When designing efficient algorithms, one also wishes to optimize memory usage and message sizes, as well as running times. Unfortunately, in the worst case, a view of Yamashita–Kameda truncated at depth $t$ grows exponentially in $t$, making their approach unsuitable for systems where memory limitation is an issue.
For this reason, Tani proposed a method to efficiently compress a truncated view by identifying isomorphic subtrees within the view itself. Using Tani’s technique, a compressed view (of an undirected network with port numbers) truncated at depth $t$ can be stored in $O(tnM\log M)$ bits, where $M$ is the maximum degree of an agent in the network [63].
This essentially matches the upper bound on the size of a vista of a history tree of a $\tau$-union-connected network at round $t$, which is $O(tn^{2}\log\tau M)$ bits (the small discrepancy is due to the fact that our networks are modeled by disconnected dynamic multigraphs, as opposed to connected static simple graphs). Thus, for instance, our linear-time terminating Counting algorithm based on history trees requires $O(n^{3}\log\tau M)$ bits of memory per agent in the worst case.
@@
A similar notion of causal influence can be formulated within the framework of history trees by stating that an agent $p$ exerts a causal influence on an agent $q$ at round $t$ if the vista of $q$ at round $t$ includes a node $v$ representing $p$. It is important to note that not all agents represented by $v$ necessarily initiate a sequence of messages reaching $q$ by round $t$. However, at least one agent represented by $v$ does so, and the specific identity of this agent is irrelevant, since agents are anonymous.
@@
Kuhn et al. established that, in a 1-union-connected dynamic network, the set of agents that are causally influenced by a given agent grows by at least one unit at every round, until all $n$ agents have been influenced. We can rephrase this observation in the language of history trees as follows.
  • The vista of any agent at any round $t$ contains nodes representing at least $\min\{t+1,n\}$ agents.
  • For any agent $p$ and any round $t$, there are at least $\min\{t+1,n\}$ agents whose vista at round $t$ contains a node representing $p$.
Broadcasting speed. As a consequence of the above, broadcasting a message, i.e., forwarding a message from a single agent to the whole dynamic network, takes at most $n-1$ rounds. Therefore, the dynamic diameter $d$, defined as the maximum number of rounds it takes to forward a message from an agent to any other agent at any point in time, satisfies $d\leq n-1$ in any 1-union-connected dynamic network.
@@
Let $P$ be a set of agents in a 1-union-connected dynamic network of size $n$, such that $11\leq|P|\leq n-$ 1, and let $t\geq 0$. Then, at every round $t^{\prime}\geq t+|P|$, in the history of every agent there is a node at level $L_{t}$ representing at least one agent not in $P$.
@@
Let $Q$ be the complement of $P$ (note that $Q$ is not empty), and let $Q_{t+i}$ be the set of agents represented by the nodes in $L_{t+i}$ whose vista contains a node in $L_{t}$ representing at least one agent in $Q$. We will prove by induction that $|Q_{t+i}|\geq|Q|+i$ for all $0\leq i\leq|P|$. The base case holds because $Q_{t}=Q$. The induction step is implied by $Q_{t+i}\subsetneq Q_{t+i+1}$, which holds for all $0\leq i<|P|$ as long as $|Q_{t+i}|<n$. Indeed, because $G_{t+i+1}$ is connected, it must contain a link between an agent $p\in Q_{t+i}$ and an agent $q\notin Q_{t+i}$. Thus, the history of $q$ at round $t+i+1$ contains the history of $p$ at round $t+i$, and so $Q_{t+i}\subsetneq Q_{t+i}\cup\{q\}\subseteq Q_{t+i+1}$.
Now, plugging $i:=|P|$, we get $|Q_{t+|P|}|=n$. In other words, the history of each node in $L_{t+|P|}$ (and hence in subsequent levels) contains a node at level $L_{t}$ representing an agent in $Q$. ∎